OpenStudy (anonymous):

Mr.Brown wants to make a rectangular pen for animals. He plans to use at least 120 feet of fencing,but no more than 200 feet. He wants the length of the pen to be three times the width and the width to be at least 30 feet. Use an inequality to show that Mr.Brown cannot do what he plans to do

OpenStudy (anonymous):

so we have \[120 \le \] feet of fencing \[\le 200\] And we know that feet of fencing = 2(3*w) + 2(w) \[w \ge 30\] So try the feet of fencing equation with the smallest possible width (which would be 30): feet of fencing= 2(3*30) +2(30)= 240 since 240 ft. is greater then the maximum amount of fencing (200) it is impossible